# Towards Cooperative Maneuver Planning in Mixed Traffic at Urban Intersections

### v1

@@ Abstract
Connected automated driving promises a significant improvement of traffic efficiency and safety on highways and in urban areas. Apart from sharing of awareness and perception information over wireless communication links, cooperative maneuver planning may facilitate active guidance of connected automated vehicles at urban intersections. Research in automatic intersection management put forth a large body of works that mostly employ rule-based or optimization-based approaches primarily in fully automated simulated environments. In this work, we present two cooperative planning approaches that are capable of handling mixed traffic, i.e., the road being shared by automated vehicles and regular vehicles driven by humans. Firstly, we propose an optimization-based planner trained on real driving data that cyclically selects the most efficient out of multiple predicted coordinated maneuvers. Additionally, we present a cooperative planning approach based on graph-based reinforcement learning, which conquers the lack of ground truth data for cooperative maneuvers. We present evaluation results of both cooperative planners in high-fidelity simulation and real-world traffic. Simulative experiments in fully automated traffic and mixed traffic show that cooperative maneuver planning leads to less delay due to interaction and a reduced number of stops. In real-world experiments with three prototype connected automated vehicles in public traffic, both planners demonstrate their ability to perform efficient cooperative maneuvers.
Index Terms:
Connected automated driving, cooperative planning, driver model, scenario prediction, optimization, reinforcement learning, graph neural network, mixed traffic, simulation, real-world evaluation, traffic efficiency.6.8in(0.85in,0.15in)
@@
I Introduction
@@
Urban traffic is prone to inefficiencies and disturbances due to the ever increasing volume of traffic. This manifests, e.g., at smaller intersections, where static priority rules are the prevailing method for coordination of vehicles. Occlusions through buildings or other road users limit the field of view for both human drivers and vehicle-bound sensors of an automated vehicle (AV). Connected automated driving opens up new opportunities to improve urban traffic efficiency by leveraging communication links between vehicles and possibly infrastructure systems. Moreover, edge computing resources are becoming available in urban areas that enable, e.g., populating and maintaining a collective environment model (EM) of an area of interest. This setup allows connected automated vehicles (CAVs) to improve their local planning algorithms by incorporating data from the server-side EM, as it was shown in [1] at a suburban three-way intersection in Ulm-Lehr, Germany.
@@
In the current work, we build upon this approach and investigate the potential of cooperative maneuver planning, as illustrated in Fig. 1. By planning such maneuvers on a centralized instance, like an edge server, CAVs can be instructed to explicitly deviate from priority rules to reach peak efficiency. Urban automotive traffic will not be fully automated in the near future. Instead, mixed traffic, i.e., the simultaneous use of roads by both human-driven vehicles (HDVs) and AVs, will be prevalent. Therefore, any cooperative planner must consider HDVs that cannot be directly influenced and behave according to the priority rules. We present two approaches to cooperative maneuver planning, one of which leverages a fast long-term multi-scenario prediction to derive an optimal maneuver, based on [2]. The second approach employs reinforcement learning (RL) to train a graph neural network (GNN) policy for cooperative maneuver planning [3]. The execution of the cooperative maneuvers in both cases is based on the coordination protocol proposed in [4]. In principle, also connected vehicles (CVs) that are driven by a human can be included in the maneuver coordination scheme. Such an application is not in focus for this paper, though. The core contribution of the present work is twofold:
@@
Statistical analysis and comparison of both approaches in a new holistic simulation framework for fully automated and mixed traffic;
@@
Demonstration of the real-world applicability of both approaches using three prototypical CAVs.
@@
The remainder of the paper is structured as follows: Section II discusses the state of the art in classical and learning-based cooperative planning for automated driving. The proposed optimization-based maneuver planner is introduced in Sec. IV, followed by the RL-based one in Sec. V. Afterwards, we present simulative evaluation results (Sec. VI) and our real-world experiments in public traffic (Sec. VII). Finally, Sec. VIII summarizes the results of the article and gives an outlook on future work.
II Related Work
@@
The efficient coordination of CAVs at urban intersections receives strong research interest in the field of automatic intersection management (AIM). Prior works mostly consider fully automated traffic and simulative evaluations, as surveyed by [5]. In the following, we give an overview on published works that are deemed most relevant to cooperative maneuver planning in mixed traffic.
Static priority rules at an intersection can be insufficient to efficiently handle a higher amount of incoming traffic and lead to long queues. Instead, usually traffic lights are used as a basic means of traffic control. The traffic light switching plan can either be determined offline, potentially optimized for the specific intersection as in [6], or adapted dynamically with a Markovian approach as in [7]. With an increasing prevalence of connected vehicles, however, the traffic at intersections can be governed by an automatic intersection management. With a fine-grained control by controlling individual vehicles, the traffic efficiency can potentially be increased compared to simpler solutions.
Many authors focus on a simpler scenario with a full deployment of only CAVs. The enumeration of all possible vehicle crossing orders as in [8] yields to the optimal ordering. However, the number of order combinations is exponential in the number of vehicles as shown in [9]. Therefore, that method is not feasible in a real-time system and a more efficient algorithm needs to evaluate only the most promising combinations. A straightforward approach is to use a heuristics to directly generate a single crossing order. One of the earliest proposed heuristics is a first-come, first-served scheme to reserve cells in a discretized space-time grid for each vehicle, which was used by Dresner and Stone in [10], by Wuthishuwong and Trächtler in [11], and others. Another heuristics-like decentralized solution for two crossing vehicles is presented in [12]. In that work, the planned vehicle trajectories are compared and the intersection management logic only intervenes as little as possible to prevent a collision.
On the other hand, some approaches try to find a near-optimal crossing order of the arriving vehicles using more elaborated algorithms. Possible optimizing strategies are dynamic programming [13], a control policy based on Petri Nets [14], or Mixed Integer Quadratic Programs [15]. In some works, the tree structure of the chronological right of way decisions between vehicles is exploited and the optimal solution is searched using branch and bound [16] or Monte Carlo Tree Search [17].
Although non-learning approaches are still prevalent in AIM, machine learning is also gaining traction in connected automated driving. In [18], it is proposed to leverage collective perception to improve the performance of local planning algorithms. The planning task only considers one AV without cooperative objective, though. Due to lack of ground-truth data for training, learning-based approaches to AIM typically rely on RL. Such a learning-based approach is proposed by [19] that suggests to train a policy through RL to choose from a restricted action space that ensures collision-free maneuvers. Cooperative maneuvers can only be initiated implicitly by vehicle behavior because there is no explicit communication between the involved vehicles. In our previous works [20, 21], we have proposed a flexible graph-based scene representation and an RL training scheme for AIM in fully automated traffic. It was shown that the model outperforms a FIFO baseline and generalizes within certain limits to intersection layouts not encountered during training.
@@
Fewer researchers have tackled the intersection management task for much more challenging mixed traffic and especially scenarios with a vast majority of HDVs, which will be prevalent in the near future. Dresner and Stone extended their cell reservation system from previous works by different fallback modes in [22]. Those modes use dynamically controlled traffic light phases to allow safe crossing of HDVs, which still provides comparable efficiency results at a rate of HDVs. Bento et al. employ a similar system that switches traffic lights to green for individual vehicles [23]. Their system reserves cells for HDVs on all possible turning lanes assuming a velocity interval to predict the maximum intersection occupancy intervals, which works well for a low percentage of HDVs. Commonly, the gains of approaches for mostly CAV traffic deteriorate in an environment with a lot of HDVs.
The approach proposed in our previous work in [2], which we extended for this paper, is specifically designed to also work with a high HDV rate. It employs a multi-scenario prediction represented in a probabilistic decision tree and optimizes the expected efficiency by maneuvers between the present CAVs. Thereby, already a deployment rate of CAVs yields to a noticeable overall efficiency gain which also benefits the HDVs.
@@
In [24], the authors propose an RL approach assuming that automated vehicles only perceive their immediate surroundings. The RL policy is shown to manage intersection traversals in presence of simulated human-driven vehicles while turning maneuvers are disallowed. A further RL framework that combines local observations with a joint reward to accommodate a cooperative objective was presented by [25]. Although this model is capable of handling mixed traffic, the network architecture limits the maximum number of cooperating vehicles in the scene. Our RL-based cooperative planning model has been extended to mixed traffic in [3]. In addition, the integration of the trained RL policy with a sampling-based motion planner was addressed in [26]. The applicability to a real test vehicle was shown in experiments on a test track.
III Cooperative Maneuver Planning Module
@@
In our proposed system architecture, the centralized cooperative planning module on the edge server plans a joint maneuver on behavior level, which is distributed to all CAVs in the scene using the protocol proposed in [4]. Each of the CAVs runs a local motion planner to derive a viable trajectory. Typical motion planning algorithms require a specification of the desired motion for a sufficient planning horizon (at least multiple seconds) into the future. Therefore, the distributed maneuver representation shall enable the vehicles to individually plan their trajectory locally.
@@
The derived maneuvers are passed to the CAVs according to the maneuver coordination protocol proposal in [4]. While the proposed protocol is very generic and supports multiple cooperative use cases, we focus on the components that are relevant for AIM. Thus, the final step to provide the planned cooperative maneuver to the CAVs is to derive so-called longitudinal maneuver waypoints defined as
@@
IV Proposed Optimization-Based Planner
The optimization-based planner builds upon our previous works presented in [2, 29]. The main component is a scenario prediction module that estimates the behavior and velocities of the vehicles in the currently observed traffic scene in the immediate future. Multiple such predictions with various applicable maneuvers are performed and the best one—according to validity and efficiency metrics—is selected to be communicated to the connected vehicles.
@@ IV Proposed Optimization-Based Planner
IV-A Driver Model for Prediction
Vital to the maneuver planning is our scenario prediction module, which we present in more detail in [29]. We employ a longitudinal lane-based prediction module trained on real-world traffic data. The driver model comprises two multi-layer perceptrons (MLPs): estimates the acceleration during the next timestep whereas determines the gap acceptance, i.e., the decision whether a vehicle should enter the intersection before another prioritized vehicle. The driver model is evaluated jointly for all vehicles in the traffic scene and integrated over a horizon with a time step of for each evaluated scenario.
@@ IV-A Driver Model for Prediction
and are the observations of towards its environment and towards a conflicting vehicle , respectively. The values are shown in Table I and Fig. 3.
@@
IV-B Priority Assignment Search
The internal maneuver representation is a list of priority assignments between pairs of cooperative vehicles. Each entry is a pair of vehicles , where the first is prioritized over the second. The planner operates cyclically in a frequency of and tries to build upon the result from previous cycles to optimize the efficiency.
As an efficiency metric to evaluate potential maneuvers, we use the relative velocity integrated over a prediction horizon that we already introduced in [2] with an additional small penalty on the maneuver complexity:
@@ IV-B Priority Assignment Search
IV-C Derivation of the Cooperative Maneuver
@@
V Proposed Learning-Based Planner
This section introduces our proposed learning model (Sec. V-A) for cooperative planning in mixed traffic, the graph-based representation (Sec. V-B), and the reward function (Sec. V-C). Finally, Sec. V-D discusses the integration of the RL policy with dedicated motion planning algorithms. We build upon our previous works [3] and [26].
V-A Learning Model
@@
The traffic environment in the area of operation is denoted by a state . This includes the full state information on all CAVs and HDVs. The environment is transferred into the next state by applying an action from the joint action set. denotes the transition function that yields the probability of changing from to under action as . Because human reaction to a given traffic state is not deterministic, the behavior of HDVs is neither, which manifests in a nondeterministic state transition. The desired policy behavior is provided through a scalar reward signal given by the function , whose detailed setup is discussed in Sec. V-C. The cooperative planner cannot observe the entire abstract traffic state because, e.g., the maneuver intention of human drivers is unknown. A smaller observation set is defined to resemble the state information that is available to the cooperative planner through the communication link or infrastructure perception. The mapping is not injective and additionally models the influence of measurement uncertainties.
@@
V-B Graph-based Representation and Network
@@
While CAVs share their desired route with the cooperative planner, the intention of human drivers is unknown. Therefore, the cooperative planner has to consider all possible conflicts with HDVs when coordinating CAVs. The graph-based scene representation inherently supports this by adding edges to an HDV for all conflicts that cannot be ruled out reliably. In the present work, we match the vehicle’s pose to lanes in a high definition (HD) map and enumerate viable continue paths. This assignment allows an angle between the vehicle heading the lane of up to at the matched point. When crossing the intersection, a vehicle’s lane matches will thus reduce to the actual destination eventually, which then allows to plan the maneuver with less excess edges. The delay of this decision might be reduced further by employing a prediction algorithm, such as [33], which is out of scope for this work, though.
@@
V-C Reward Function Engineering

### v2

# Comparison of two Cooperative Maneuver Planning Approaches at a Real-World T-Junction

@@ Abstract
Connected automated driving promises a significant improvement of traffic efficiency and safety on highways and in urban areas.
Cooperative maneuver planning may facilitate active guidance of connected automated vehicles at intersections.
Research in automatic intersection management put forth a large body of works that mostly employ rule-based or optimization-based approaches primarily in fully automated simulated environments.
In this work, we compare two cooperative planning approaches for unsignalized intersections that are capable of handling mixed traffic, i.e., the road being shared by automated vehicles and regular vehicles driven by humans.
The first approach is a cooperative planner that selects the most efficient out of multiple possible maneuvers based on a scene prediction trained on real driving data.
The second cooperative planning approach is based on graph-based reinforcement learning, which conquers the lack of ground truth data for cooperative maneuvers.
We thoroughly evaluate both cooperative planners in a realistic high-fidelity simulation with fully automated traffic and mixed traffic.
The simulative experiments show that cooperative maneuver planning leads to less delay due to interaction and a reduced number of stops.
Furthermore, we present results from real-world experiments with three prototype automated vehicles at a T-junction in public traffic, in which both planning modules demonstrate their ability to perform efficient cooperative maneuvers.
6.89in(1.81in,1.7in)
@@
Connected automated driving, cooperative planning, driver model, scenario prediction, optimization, reinforcement learning, graph neural network, mixed traffic, simulation, real-world evaluation, traffic efficiency.
1 INTRODUCTION
@@
Urban traffic is prone to inefficiencies and disturbances due to the ever increasing volume of traffic. This manifests, e.g., at smaller intersections, where static priority rules are the prevailing method for coordination of vehicles. Connected automated driving opens up new opportunities to improve urban traffic efficiency by leveraging communication links between vehicles and possibly infrastructure systems. Moreover, edge computing resources are becoming available in urban areas that enable, e.g., populating and maintaining of a collective environment model (EM) of an area of interest. This setup allows connected automated vehicles (CAVs) to improve their local planning algorithms by incorporating data from the server-side EM, as it was shown in [1] at a suburban three-way intersection in Ulm-Lehr, Germany.
@@
Connectivity between vehicles can also be used for the coordination of maneuvers to increase traffic efficiency and safety. In the current work, we integrate and evaluate two cooperative crossing maneuver planners, as illustrated in Fig. 1. The cooperative maneuvers are planned on a centralized edge server, which instructs CAVs to explicitly deviate from priority rules to reach peak efficiency. As urban automotive traffic will not be fully automated in the near future, mixed traffic, i.e., the simultaneous use of roads by both human-driven vehicles (HDVs) and CAVs, will be prevalent. Therefore, any real-world cooperative planner must consider HDVs that cannot be directly influenced and behave according to the priority rules. We compare two approaches to cooperative maneuver planning, one of which leverages a multi-scenario prediction to derive an optimal maneuver, based on [2]. The second approach employs reinforcement learning (RL) to train a graph neural network (GNN) policy for cooperative maneuver planning [3]. To provide realistic results, we employ the real-time simulation framework DeepSIL [4] with a state-of-the-art human driver model for HDVs. We extended the simulation framework to support multiple CAVs and integrated both maneuver planners as well as a trajectory planner and controller combination [5] for CAVs, the latter of which is also used in the real-world experiments. The execution of the cooperative maneuvers in simulation and in real-world is based on the coordination protocol proposed in [6]. The core contribution of the present work is twofold:
@@
Statistical analysis and evaluation of both planning approaches in a holistic simulation framework for fully automated and mixed traffic with realistic HDV and CAV models;
@@
Demonstration of the real-world applicability of both planners using three prototypical CAVs in public traffic.
@@
The remainder of the paper is structured as follows: Section 2 discusses the state of the art in classical and learning-based cooperative planning for automated driving. The optimization-based maneuver planner is presented in Sec. 4, followed by the RL-based one in Sec. 5. Afterwards, we discuss our simulative evaluation results (Sec. 6) and real-world experiments in public traffic (Sec. 7). Finally, Sec. 8 summarizes the results of the article and gives an outlook on future work.
2 RELATED WORK
@@
The efficient coordination of CAVs at urban intersections receives strong research interest in the field of automatic intersection management (AIM). In the following, we give an overview on published works that are deemed most relevant to cooperative maneuver planning at unsignalized intersections. Prior works mostly consider fully automated traffic and pure simulative evaluations, as surveyed by [7]. Also, the evaluation of the approaches often does not consider the processing time, and many researchers do not consider a communication protocol to coordinate the traffic. Therefore, most previous works are not suitable for real-world deployment.
The enumeration of all possible vehicle crossing orders in CAV-only traffic yields to the optimal ordering [8]. However, the number of order combinations is exponential in the number of vehicles as shown in [9], thus, exhaustive methods are not feasible in a real-time system. A more efficient approach is to use a heuristics to directly generate a single crossing order. One of the earliest proposed heuristics is a first-come, first-served scheme to reserve cells in a discretized space-time grid for each vehicle, which was used by several research groups [10, 11]. Other approaches try to find a near-optimal crossing order of the arriving vehicles using more elaborated algorithms. Possible optimizing strategies are dynamic programming [12], ant colonies [9], a control policy based on Petri Nets [13], Mixed Integer Quadratic Programs [14], or Monte Carlo Tree Search [15]. However, these works are designed for CAV-only traffic, and only the authors in [13] show the real-time capability of their algorithm in simulation and in real-world experiments. The other methods are tested in synthetic non-real-time simulations only, not considering realistic traffic behavior and processing delays.
@@
Fewer researchers have tackled the intersection management task for much more challenging mixed traffic and especially scenarios with a vast majority of HDVs, which will be prevalent in the near future. Previous mixed traffic approaches employ signalized intersection management to support the presence of HDVs. Dresner and Stone extended their cell reservation system from previous works by different fallback modes in [16]. Those modes use dynamically controlled traffic light phases to allow safe crossing of a low rate of HDVs (). Bento et al. employ a similar system that switches traffic lights to green for individual vehicles [17], which works well for a low percentage of HDVs. For both approaches, the gains deteriorate in an environment with a lot of HDVs. Two different research groups [18, 19] proposed virtual platoons for signalized intersection management supporting mixed traffic at high HDV rates. Their methods use a heuristics and a branch-and-bound algorithm, respectively, and are evaluated in a traffic simulation demonstrating real-time capability. However, all these methods require traffic lights to control HDVs. Also, most approaches assume that the turn direction of HDVs is known, which is generally not given in real-world traffic.
The approach proposed in our previous work in [2] is specifically designed to work at unsignalized intersections with any HDV rate and unknown HDV turn directions. It employs a scene-consistent multi-scenario prediction, which we enhanced for this paper, and optimizes the expected efficiency by maneuvers between the present CAVs. A respective coordination protocol to execute the cooperative maneuvers on CAVs was previously proposed in [6], making this approach applicable on real vehicles.
Although non-learning approaches are still prevalent in AIM, machine learning is also gaining traction in connected automated driving. In [20], it is proposed to leverage collective perception to improve the performance of local planning algorithms. The planning task only considers one automated vehicle without cooperative objective, though. Due to lack of ground-truth data for training, learning-based approaches to AIM typically rely on RL. Such a learning-based approach is proposed by [21] that suggests to train a policy through RL to choose from a restricted action space that ensures collision-free maneuvers. However, this approach is designed for CAV-only traffic and the evaluation relies solely on simulations.
@@
In [22], the authors propose an RL approach assuming that automated vehicles only perceive their immediate surroundings. The RL policy is shown to manage intersection traversals in presence of simulated human-driven vehicles while turning maneuvers are disallowed. A further RL framework that combines local observations with a joint reward to accommodate a cooperative objective was presented by [23]. Although this model is capable of handling mixed traffic, the network architecture limits the maximum number of cooperating vehicles in the scene. These works also lack the application to real vehicles and a comprehensive evaluation in public traffic.
In our previous works [24, 25], we have proposed a flexible graph-based scene representation and an RL training scheme for AIM in fully automated traffic. It was shown that the model outperforms a FIFO baseline and generalizes within certain limits to intersection layouts not encountered during training. Our RL-based cooperative planning model has been extended to mixed traffic in [3]. This approach also does not rely on the turn direction of the HDVs to be known. In addition, the integration of the trained RL policy with a sampling-based motion planner was addressed in [26], demonstrating the applicability on real vehicles.
3 COOPERATIVE MANEUVER PLANNING MODULES
@@
In our proposed system architecture, the centralized cooperative planning module on the edge server plans a joint maneuver on behavior level, which is distributed to all CAVs in the scene. Each of the CAVs runs a local motion planner to derive a viable trajectory. Typical motion planning algorithms require a specification of the desired motion for a sufficient planning horizon (at least multiple seconds) into the future. Therefore, the distributed maneuver representation shall enable the vehicles to individually plan their trajectory locally.
@@
The derived maneuvers are passed to the CAVs according to the maneuver coordination protocol proposal in [6]. While the protocol is very generic and supports multiple cooperative use cases, we focus on the components that are relevant for AIM. Thus, the final step to provide the planned cooperative maneuver to the CAVs is to derive so-called longitudinal maneuver waypoints defined as
@@
4 OPTIMIZATION-BASED PLANNER
The optimization-based planner builds upon our previous works presented in [2]. The main component is our scene-consistent prediction module [29] that estimates the behavior and velocities of the vehicles in the currently observed traffic scene in the immediate future. Multiple such scenario predictions with various applicable maneuvers are performed and the best one—according to validity and efficiency metrics—is selected to be communicated to the connected vehicles.
@@ 4 OPTIMIZATION-BASED PLANNER
4.1 Driver Model for Prediction
Vital to the maneuver planning is the scenario prediction module. We employ a longitudinal lane-based prediction module trained on real-world traffic data. The driver model comprises two multi-layer perceptrons (MLPs): estimates the acceleration during the next timestep whereas determines the gap acceptance, i.e., the decision whether a vehicle should enter the intersection before another prioritized vehicle. The driver model is evaluated jointly for all vehicles in the traffic scene and integrated over a horizon with a time step of for each evaluated scenario.
@@ 4.1 Driver Model for Prediction
and are the observations of towards its environment and towards a conflicting vehicle , respectively. The values are shown in Table 1 and Fig. 3.
@@
4.2 Priority Assignment Search
The internal maneuver representation is a list of priority assignments between pairs of cooperative vehicles. Each entry is a pair of vehicles , where the first is prioritized over the second. The planner operates cyclically in a frequency of and tries to add assignments to the resulted from previous cycles to optimize the efficiency.
As an efficiency metric to evaluate potential maneuvers, we use the relative velocity integrated over a prediction horizon with an additional small penalty on the maneuver complexity:
@@ 4.2 Priority Assignment Search
4.3 Derivation of the Cooperative Maneuver
@@
5 LEARNING-BASED PLANNER
This section introduces our proposed learning model (Sec. 5.1) for cooperative planning in mixed traffic, the graph-based representation (Sec. 5.2), and the reward function (Sec. 5.3). Finally, Sec. 5.4 discusses the integration of the RL policy with dedicated motion planning algorithms. We combine the methods presented in our previous works [3] and [26] and provide a brief summary of the main features.
5.1 Learning Model
@@
The traffic environment in the area of operation is denoted by a state . This includes the full state information on all CAVs and HDVs. The environment is transferred into the next state by applying an action from the joint action set. denotes the transition function that yields the probability of changing from to under action as . Because human reaction to a given traffic state is not deterministic, the behavior of HDVs is neither, which manifests in a nondeterministic state transition. The desired policy behavior is provided through a scalar reward signal given by the function , whose detailed setup is discussed in Sec. 5.3. The cooperative planner cannot observe the entire abstract traffic state because, e.g., the maneuver intention of human drivers is unknown. A smaller observation set is defined to resemble the state information that is available to the cooperative planner through the communication link or infrastructure perception. The mapping is not injective and additionally models the influence of measurement uncertainties.
@@
5.2 Graph-based Representation and Network
@@
While CAVs share their desired route with the cooperative planner, the turn direction of human drivers is unknown. Therefore, the cooperative planner has to consider all possible conflicts with HDVs when coordinating CAVs. The graph-based scene representation inherently supports this by adding edges to an HDV for all conflicts that cannot be ruled out reliably. In the present work, we match the vehicle’s pose to lanes in a high definition (HD) map and enumerate viable continue paths. This assignment allows an angle between the vehicle heading the lane of up to at the matched point. When crossing the intersection, a vehicle’s lane matches will thus reduce to the actual destination eventually, which then allows to plan the maneuver with less excess edges. The delay of this decision might be reduced further by employing a prediction algorithm, such as [33], which is out of scope for this work, though.
@@
5.3 Reward Function Engineering